### v2

# MATHWELL: Generating Educational Math Word Problems at Scale

Abstract
Math word problems are critical K-8 educational tools, but writing them is time-consuming and requires domain expertise. We suggest that language models can support K-8 math education by automatically generating problems at scale. To be educational, generated problems must be 1) solvable, 2) accurate, and 3) appropriate. Existing datasets are unlabeled for these criteria, making them ill-suited for training problem generators. We introduce MATHWELL, a Llama-2 (70B) model iteratively finetuned to generate K-8 math word problems using data from expert annotation. Using MATHWELL, we generate the largest English word problem dataset with Program of Thought (PoT) rationales to date, containing 20,490 problems. 3,484 are scored by domain experts who find MATHWELL has a 40% higher share of problems that have executable solutions and meet all criteria than alternatives, with 74% of its problems with executable solutions being solvable, accurate, and appropriate. We release our model, data, and annotations.111https://github.com/bryanchrist/MATHWELL
1 Introduction
Math word problems are a critical part of assessing student learning in K-12 education Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). Customizing word problems to student interests can further increase learning, interest in math, and test performance Bernacki and Walkington (2018); Walkington (2013). However, with many students and responsibilities, teachers rarely have time to customize questions for their students. We aim to automatically generate customized math word problems and answers at scale to facilitate fast, interest-guided, math education.
Math word problems include natural language questions paired with numerical solutions, as illustrated in Figure 1. To be effective educational tools for grade-school students, the problems must be solvable, appropriate, and have accurate solutions. While most public word problem resources are manually curated and, therefore, quite limited (e.g., K5Learning ; TPT ; VDOE ), there have been recent efforts to generate math word problems using natural language processing Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Norberg et al. (2023); Qin et al. (2023); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023). Closest to our work, some recent efforts explore math word problem generation using Large Language Models (LLMs) Niyarepola et al. (2022); Zong and Krishnamachari (2023). However, to the best of our knowledge, existing methods have one or both of two important limitations. First, many only generate questions (e.g., Niyarepola et al. 2022; Zong and Krishnamachari 2023). To be educationally useful, students and teachers must also have access to answers. Further, it is hard to verify if questions are solvable and appropriate without answers. Second, all existing methods are context-dependent, simply rephrasing or creating new output based on input problems or equations. In practice, curating and choosing specific word problems or equations beforehand is laborious and difficult Roche (2013).
To address the aforementioned limitations, we propose context-free educational math word problem generation. Successful solutions must return questions and answers together and require no hand-picked equations or exemplar word problems. For context-free word problem generators to be effective educational tools, we propose three criteria to evaluate their outputs. 1) Each problem must be solvable. 2) Each question’s corresponding answer must be accurate. 3) Each problem must be appropriate, where the math topic should be familiar to a student and the question’s context should be appropriate for a young learner.
We propose Math Assessment Through High-interest Word Problems Extracted from LLMs (MATHWELL), the first context-free educational math word problem generator. MATHWELL is a Llama-2 (70B) model finetuned on popular math question-answering (QA) datasets used by recent math LLMs Azerbayev et al. (2023); Luo et al. (2023); Yue et al. (2023). However, these datasets are unannotated for solvability, accuracy, and appropriateness and are not designed to train word problem generators. We therefore conduct multiple rounds of annotation by real expert educators, which we use to progressively finetune MATHWELL. With these annotations, our resultant model’s outputs are better targeted for grade-school math education than outputs from alternative general-purpose math LLMs. Further, we train MATHWELL to generate Program of Thought (PoT) solutions to word problems, which recent methods show improves accuracy on open-ended questions Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023).
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems with PoT rationales. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM has a more age appropriate reading level than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs.
Our work has the following contributions:
We introduce a novel context-free word problem generation task and, through expert human annotation, propose solvability, accuracy and appropriateness as the primary metrics for evaluating output from these models.
We release MATHWELL, a context-free word problem generator, and show it has a 43.4% higher share of outputs that meet all evaluation criteria and have executable code than the next best open-source model and is the only model that generates questions that are all at an appropriate reading level for K-8 students.
We release SGSM and show it is high quality through automatic evaluation. SGSM is the only existing dataset designed to train context-free word problem generators, is the largest English math grade school dataset with PoT solutions, and contains more readable questions than existing grade school math datasets.
2 Related Work
Mathematical Reasoning in LLMs
Mathematical reasoning as measured by performance on math QA benchmarks such as MATH Hendrycks et al. (2021b), MMLU-STEM Hendrycks et al. (2021a), and GSM8K Cobbe et al. (2021) is one critical way LLMs are evaluated, and much recent work has focused on how to improve mathematical problem solving in LLMs. Common approaches include domain-specific finetuning Azerbayev et al. (2023); Lewkowycz et al. (2022), prompting techniques Gao et al. (2023); Wei et al. (2023), external solution verifiers Cobbe et al. (2021), and majority voting among candidate solutions Wang et al. (2023). Because we use existing math QA datasets to train our model to generate questions/answers simultaneously, we use prompting techniques to improve mathematical reasoning. We prompt our model to generate Program of Thought (PoT) solutions rather than Chain of Thought (CoT) solutions because recent work has found that PoT outperforms CoT for open-ended questions, and all K-8 word problems are open ended Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023).
Math QA Datasets
There are several existing datasets for training LLMs to answer grade school math questions in English, including GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022b), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). However, because these datasets were created to train LLMs to answer, rather than generate, grade school math problems, they do not possess the characteristics we consider critical to training LLMs to generate solvable, accurate, and appropriate question/answer pairs. Specifically, while some datasets include PoT solutions (GSM-Hard, MathInstruct GSM8K, and NumGLUE), only GSM-Hard contains Python function solutions, which we find increases the likelihood that a question/answer pair contains executable code. Because GSM-Hard contains questions with inflated number values, though, it is not an appropriate dataset for training a question generator for K-8 students. Due to the limitations of existing datasets, we create and release SGSM, a SGSM subset verified by human annotation that contains Python function solutions and appropriate number ranges for K-8 students.
Math Word Problem Generation
Other work explores automatically generating math word problems but takes a more limited approach than the present study. The closest related work is Zong and Krishnamachari (2023), who assess GPT-3’s ability to generate math word problems. Their method is not context-free, however, as they use a reference problem to guide generation. In addition, they only generate word problems, not solutions. Niyarepola et al. (2022) train LLM word problem generators, but they guide generation by providing the beginning of a word problem and do not generate solutions. Most other existing works use LLMs or deep neural networks to generate word problems based on pre-specified equations, provide additional questions based on reference word problems, or re-write existing word problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Norberg et al. (2023); Qin et al. (2023); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023), each of which are more limited than context-free generation by restricting the range of possible outputs. These approaches also require additional input from users, which may not be feasible for teachers or students who wish to create customized word problems. As a result, we do not compare MATHWELL against these models/methods. Our study is the first to evaluate the capability of leading open-source LLMs for the task of math word problem generation.
3 Methods
We propose MATHWELL, a finetuned language model that generates word problems that align with topics K-8 students are interested in. MATHWELL can be prompted to generate word problems, where each problem is a question/answer pair. Then, we use this language model to generate a large synthetic dataset, which we evaluate with human domain experts. We illustrate our model finetuning and dataset generation process in Figure 2.
There are three critical criteria for model-generated question/answer pairs: 1) Solvability indicates questions are possible to solve and have one correct answer Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Qin et al. (2023). 2) Accuracy measures if the generated answer is correct. 3) Appropriateness indicates that questions make logical sense, do not have grammatical errors or conflicting information, and are appropriate for K-8 students mathematically and contextually. A secondary characteristic we consider for model comparisons is topic specificity, or if the question incorporates the topic specified in the prompt.
| Dataset | N | PoT | PF | AD | AL | FKGL | NDC | BF1 |
| GSM-Hard Gao et al. (2023) | 1,319 | ✗ | 72.9 (25.6) | 4.21 (2.43) | 8.20 (1.13) | 84.0 | ||
| MathInstruct GSM8K Yue et al. (2023) | 6,403 | ✗ | 66.2 (23.9) | 4.25 (2.48) | 8.17 (1.13) | 84.6 | ||
| NumGLUE Mishra et al. (2022a) | 12,403 | ✗ | ✗ | 144.8 (136.5) | 10.04 (6.99) | 10.27 (1.51) | 81.5 | |
| ASDIV Miao et al. (2020) | 2,305 | ✗ | ✗ | 45.1 (15.8) | 3.56 (2.40) | 7.85 (1.48) | 85.5 | |
| SVAMP Patel et al. (2021) | 1,000 | ✗ | ✗ | 47.3 (11.7) | 3.39 (2.07) | 7.84 (1.09) | 86.1 | |
| SGSM (Ours) | 20,490 | ? | 62.0 (15.0) | 2.68 (1.97) | 7.99 (1.26) | 84.8 | ||
| SGSM | 2,093 | 57.2 (15.7) | 2.50 (1.76) | 8.12 (1.25) | 85.2 | |||
| SGSM | 18,397 | ? | 62.5 (14.8) | 2.70 (1.99) | 7.97 (1.26) | 84.9 |
3.1 Training MATHWELL
MATHWELL is a finetuned Llama-2 (70B) Touvron et al. (2023) model that generates question/answer pairs. After initial finetuning using public data, we generate synthetic data and acquire expert annotations for solvability, accuracy, and appropriateness. We then further finetune on the best samples to ensure high-quality outputs. We generate Program of Thought (PoT) solutions, which recent works have shown improves accuracy Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023).
Initial Finetuning
We re-format MathInstruct GSM8K Yue et al. (2023) into an Alpaca-style Taori et al. (2023) instruction dataset of question/answer pairs with a standard prompt asking the model to generate a grade school math word problem. To the best of our knowledge, MathInstruct GSM8K is the only dataset with Python code solutions and questions that are all appropriate for K-8 students Yue et al. (2023). We also include MathInstruct MATH and TheoremQA PoT Yue et al. (2023) question/answer pairs with a standard prompt asking the model to generate a challenge math problem to expose the model to more complex code and further promote mathematical reasoning. We then finetune Llama-2 (70B) on this dataset of 25,926 rows using QLoRA Dettmers et al. (2023) for 4,250 steps. We use two A100 GPUs, a learning rate of 1e-4, LoRA modules at every model layer, per-device batch size of 1, and a 3% warm-up ratio.
Secondary Finetuning on Expert-annotated Data
Inspired by recent works that iteratively finetune LLMs Guo et al. (2024); Wang et al. (2024), including OpenAI’s FeedME approach Zheng et al. (2023), we next aim to evaluate and improve how well MATHWELL meets our three main criteria: solvability, accuracy, and appropriateness. However, these properties are unannotated in prior works. Therefore, we few-shot prompt the finetuned model and generate 3,234 problems. We then acquire annotations for solvability, accuracy, and appropriateness from volunteers with previous K-12 teaching experience/training. Further details about the annotation process and annotator agreement are in Appendix F. These annotations identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL on these MaC outputs for 1,250 more steps.
3.2 Generating a Dataset with MATHWELL
We next prompt MATHWELL to generate a large dataset, which we call Synthetic Grade School Math (SGSM). SGSM contains 20,490 question/answer pairs generated by MATHWELL. SGSM has two subsets: SGSM, containing 2,093 MaC question/answer pairs, and SGSM, containing 18,397 question/answer pairs that have executable code but are not labeled for solvability, accuracy, and appropriateness. SGSM is the largest English grade school PoT math QA dataset. To support research in automatically scoring model-generated math word problems, we release our annotated data. To the best of our knowledge, this is the only human annotated model-generated math word problem dataset (see Appendix B for details).
| Model | Solv. | Acc. | App. | MaC | Top. Spec. | EC | EC/MaC |
| LLEMMA | 48.8 (3.17) | 63.9 (4.37) | 41.8 (4.48) | 15.2 (2.28) | 94.8 (1.41) | 24.3 (0.70) | 3.70 (0.55) |
| MAmmoTH | 86.8 (2.15) | 94.9 (1.49) | 67.7 (3.18) | 56.8 (3.14) | 97.6 (0.97) | 6.90 (0.36) | 3.91 (0.22) |
| Llama-2 | 84.0 (2.32) | 89.5 (2.12) | 81.0 (2.72) | 62.4 (3.07) | 99.2 (0.56) | 55.4 (0.98) | 34.6 (1.70) |
| MATHWELL | 89.2 (1.97) | 96.9 (1.17) | 86.5 (2.29) | 74.8* (2.75) | 99.6 (0.40) | 66.4* (1.00) | 49.6* (1.83) |
SGSM Dataset Characteristics
Table 1 shows the advantages of SGSM over existing datasets. First, it is the largest available PoT dataset, which may enable further training of task-specific models. Second, SGSM is the only dataset with both Program of Thought (PoT) solutions written as Python functions and questions that assess mathematical concepts appropriate for K-8 students, which we find critical for training context-free word problem generators (see Appendix A). Based on the experiments reported in Section 4, we believe SGSM is also predominately comprised of questions that are mathematically appropriate for K-8 students. Third, SGSM and its subsets have the lowest average Flesch-Kincaid grade level (FKGL) Flesch (1948); Kincaid et al. (1975) and comparable New Dale-Chall (NDC) readability score Chall and Dale (1995), so the questions may be more appropriate for those who struggle to read.
SGSM is similar to existing datasets on other metrics commonly used to evaluate math datasets and/or model-generated word problems (e.g., Jiao et al. 2023; Niyarepola et al. 2022; Norberg et al. 2023; Zhou and Huang 2019; Zhou et al. 2023). Specifically, as shown in Table 1, while SGSM has a shorter average token length than MathInstruct GSM8K Yue et al. (2023), its token length is longer than ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021), two other grade school math datasets. This suggests that SGSM’s question complexity, as reflected in average question length, should be similar to existing datasets. SGSM’s BERTScore is close to those of existing datasets, suggesting the questions are similar. Taken together, these findings suggest that SGSM has important advantages over existing datasets and is comparable to human-written data.
4 Experiments
We evaluate MATHWELL’s outputs using both human evaluations with domain experts and automatic evaluations. We compare 250 randomly-selected MATHWELL outputs to those of LLEMMA (34B) Azerbayev et al. (2023), MAmmoTH (70B) Yue et al. (2023), and Llama-2 (70B) Touvron et al. (2023), which are recent and popular. LLEMMA and MAmmoTH represent two different approaches to training math LLMs, as MAmmoTH is trained as a QA model Yue et al. (2023), while LLEMMA is trained to be more general using a large corpus of mathematical content Azerbayev et al. (2023). Llama-2 serves as a baseline to ensure task-specific finetuning increases performance. Each model is prompted using example question/answer pairs from SGSM and asked to create a question/answer pair on a random topic K-8 students are interested in. Our 250 evaluation sample for each model is roughly five times larger than the human evaluation samples used in other math word problem generator studies Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023).
4.1 Human Evaluation
| Solvable Questions | MaC Questions | |||||||||||||
| Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. | No Ops | Add. | Sub. | Mult. | Div. | Frac. | Dec. | Total Ops |
| LLEMMA | 34.4 | 27.0 | 33.6 | 20.5 | 6.56 | 15.6 | 15.6 | 36.8 | 39.5 | 31.6 | 15.8 | 2.63 | 13.2 | 1.39 |
| MAmmoTH | 39.6 | 37.8 | 43.8 | 19.4 | 3.69 | 10.6 | 2.30 | 43.0 | 42.2 | 40.8 | 16.9 | 4.93 | 9.86 | 1.58 |
| Llama-2 | 57.6 | 58.6 | 22.9 | 14.3 | 8.10 | 11.4 | 4.76 | 59.6 | 60.3 | 24.4 | 12.8 | 5.77 | 8.97 | 1.72 |
| MATHWELL | 69.5 | 69.1 | 24.7 | 10.3 | 5.38 | 7.62 | 1.35 | 71.1 | 70.6 | 24.6 | 8.56 | 4.81 | 7.49 | 1.87 |
MATHWELL Outperforms Existing Models
Our expert annotators scored the 250 outputs from each model for solvability, accuracy, appropriateness, and topic specificity. As shown in Table 2, MATHWELL performs best in each metric, with a 19.9% higher share of outputs with executable code that MaC, 19.9% higher share that have executable code, and 43.4% higher share that have executable code and MaC than the next best model. Table 2 highlights that existing math models are not as capable at this task as Llama-2 and MATHWELL. Specifically, math models struggle with appropriateness and generating executable code, showing that raw coding capability does not necessarily translate into more outputs containing executable code. Table 2 also suggests that QLoRA finetuning on high-quality data is enough to match or exceed math models’ accuracy on this task. Although MATHWELL performs best in topic specificity, the other models do well in this metric, implying that MATHWELL and LLMs can effectively customize word problems to student interests.
MATHWELL Generates High-quality and Complex Questions
Word problems involving multiplication, division, fractions, and decimals are more complex than those only involving addition and subtraction. In Table 3, we assess whether each model can generate MaC questions when using more complicated operations. Questions may contain more than one operation/topic. Solvable questions may require no math operation if they contain the answer in the question (see Appendix E.4.3), so we also report the share of questions containing no operations. To determine if questions with complex operations are accurate and appropriate, we compare the math operations/topics in solvable questions to those in MaC questions. We also report the average number of distinct operations/topics in MaC questions for each model, which is another way to assess question complexity.
Table 3 shows that MATHWELL is the only model for which the share of MaC questions for each math operation/topic is within two percentage points of that for solvable questions, providing evidence that MATHWELL can generate MaC questions regardless of the complexity of the operation. MATHWELL is also the least likely to generate problems that require no operations and has the highest average total operations, two other pieces of evidence that suggest MATHWELL generates high-quality, complex problems.
In our evaluation, we do not prompt models for specific operations. Under these conditions, MATHWELL generates more problems containing addition and subtraction relative to the other models. In turn, there is a concern that MATHWELL’s performance in Table 2 could be due to it generating simple questions for this experiment, which may be more likely to MaC. To address this concern, we conduct two additional analyses reported in Appendix C: 1) logistic regressions showing MATHWELL’s higher MaC relative to the other models holds when controlling for math operations and 2) a summary of accuracy by operation showing MATHWELL is the only model for which accuracy does not substantially differ by operation and remains above 90% for each operation.
4.2 Automatic Evaluation
| Dataset | PPL | BF1 |
| MathInstruct GSM8K | 3.07 (0.691) | 84.6 |
| SGSM | 2.44 (0.439) | 84.3 |
| SGSM | 2.32 (0.666) | 83.9 |
| Model | PPL | BF1 | GSM BF1 | MaC AL | NDC |
| LLEMMA | 3.82 (0.10) | 84.3 | 84.1 | 50.9 (2.89) | 8.41 (0.09) |
| MAmmoTH | 2.76 (0.03) | 86.0 | 84.7 | 44.4 (1.15) | 8.25 (0.08) |
| Llama-2 | 2.52 (0.03) | 85.5 | 84.3 | 49.8 (1.19) | 8.20 (0.07) |
| MATHWELL | 2.44 (0.03) | 85.5 | 84.2 | 54.1 (0.97) | 8.23 (0.08) |
MATHWELL Outputs Have Low Perplexity
Perplexity (PPL) is commonly used to measure the quality of LLM outputs and has been used by prior works Jiao et al. (2023). Lower values imply better outputs and we calculate PPL using Llama-2 (70B). Table 4 shows SGSM has lower PPL than MathInstruct GSM8K and Table 5 shows MATHWELL’s 250 evaluation sample for the experiments in Table 2 has lower PPL than that of the other models, suggesting our synthetic data are high quality.
| Solvability | Accuracy | Appropriateness | ||||||||||||||||
| Source | Acc. | BA | P | R | F1 | RA | Acc. | BA | P | R | F1 | RA | Acc. | BA | P | R | F1 | RA |
| Annotated Test | 78.9 | 61.9 | 86.7 | 88.0 | 87.4 | 0.701 | 81.6 | 53.8 | 87.2 | 92.2 | 89.6 | 0.649 | 82.3 | 63.1 | 86.0 | 93.5 | 89.6 | 0.693 |
| LLEMMA | 48.8 | 49.8 | 48.7 | 90.2 | 63.2 | 0.542 | 59.0 | 47.1 | 62.5 | 89.7 | 73.7 | 0.492 | 43.4 | 50.0 | 41.8 | 90.2 | 57.1 | 0.450 |
| MAmmoTH | 84.4 | 52.5 | 87.4 | 95.9 | 91.4 | 0.700 | 93.5 | 57.9 | 95.7 | 97.6 | 96.6 | 0.669 | 69.1 | 54.0 | 69.6 | 96.6 | 80.9 | 0.529 |
| Llama-2 | 82.4 | 53.1 | 84.9 | 96.2 | 90.2 | 0.619 | 85.2 | 51.6 | 89.8 | 94.1 | 91.9 | 0.677 | 77.6 | 53.7 | 82.2 | 92.4 | 87.0 | 0.579 |
| MATHWELL | 85.2 | 54.3 | 90.1 | 93.7 | 91.9 | 0.599 | 94.6 | 62.7 | 97.7 | 96.8 | 97.2 | 0.790 | 83.0 | 57.8 | 88.6 | 92.2 | 90.4 | 0.711 |
MATHWELL Outputs Are Similar to Human-written Questions
Like Jiao et al. (2023) and Zhou et al. (2023), we use BERTScore Zhang et al. (2020) to compute the semantic similarity of our synthetic data and compare it to existing datasets. A lower BERTScore for synthetic questions relative to existing datasets would imply they are less similar to each other than human-written questions, while a higher score would imply they are more similar. We also use BERTScore to compare our synthetic data to MathInstruct GSM8K to identify if they are similar to human-written questions. We calculate MathInstruct GSM8K’s within-dataset BERTScore as a reference to compare each source against. As shown in Tables 4 and 5, across models, SGSM subsets and MathInstruct GSM8K comparisons, BERTScores are similar, suggesting the questions we generate and the data we release are similar to human quality.
MATHWELL Outputs Longer MaC Questions
Longer token length may signal increased question complexity, as longer word problems often contain more information and mathematical operations than shorter ones. Comparing the average length of all questions to the length of MaC questions can determine if MaC questions are shorter or simpler. As shown in Table 5, MATHWELL’s MaC questions are the longest of the models considered, suggesting its MaC problems may be more complex than those from other models. As shown in Table 9, MATHWELL is also the only model for which MaC length is within a token of the overall average, providing evidence that its MaC questions are no simpler than its average question.
MATHWELL Outputs Have Appropriate Readability
Reading level is one way to automatically assess if math word problems are written appropriately. Like Norberg et al. (2023), we use FKGL and NDC to evaluate reading level. FKGL calculates reading level as a function of the total words, total sentences, and total syllables in a piece of text, with the score representing the U.S. grade level of the text Aggarwal ; Flesch (1948); Kincaid et al. (1975). Negative FKGL scores are possible and denote text that is easy to read due to having short words and sentences. NDC calculates readability as a function of sentence length and the proportion of words not contained in a list of 3,000 common English words Aggarwal ; Chall and Dale (1995).
As shown in Tables 5 and 1, MATHWELL’s NDC is slightly higher than its training datasets but similar to the other models, remaining in the same score threshold for the metric Chall and Dale (1995). One reason for this difference could be that the models’ NDC is inflated because all questions are topic specific and include proper nouns that may be familiar to a young learner but are not in the list of 3,000 common words (e.g., Fortnite). Taken together, this evidence suggests that MATHWELL’s outputs, as measured by NDC, are likely similar to existing datasets.
Figure 3 compares the FKGL distribution of SGSM questions to the three other existing datasets that are appropriate for grade school students. Fewer of SGSM’s questions than other existing datasets are written at a grade level beyond 8th grade. This suggests that the dataset captures the intended age range and complements existing datasets by including questions more appropriate for struggling readers. Figure 4 compares the FKGL distribution of MATHWELL outputs to the other models considered. While MATHWELL has a similar average FKGL, it is the only model that does not generate questions beyond an 8th grade reading level, providing evidence it creates more readable questions for K-8 students.
4.3 Question Classifier Performance
Given the high cost of human evaluation, we consider automatically evaluating model outputs by training DistilBERT Sanh et al. (2019) classifiers for solvability, accuracy and appropriateness using our annotated data, excluding the evaluation samples from each model. We use an 80/10/10 train, validation and test split. The solvability dataset has 3,234 rows, the accuracy dataset has 2,830 rows, and the appropriateness dataset has 2,660 rows. We exclude unsolvable questions from the accuracy and appropriateness datasets to promote specialization, as these questions do not have a solution and are inappropriate. Each dataset is unbalanced, with 82.3% of questions labeled as solvable, 85.8% labeled as accurate, and 80.1% labeled as appropriate. As a result, we modify the training objective to weight the loss based on the inverse proportion of observations from each class. We train each model for 8 epochs using the HuggingFace library Wolf et al. (2020) with strict regularization (weight decay = 0.9) and use the checkpoint with the lowest validation loss for testing.
As shown in Table 6, despite weighting the loss, each classifier has a lower balanced accuracy than its overall accuracy. The classifiers perform best on output from models derived from Llama-2, although performance for all models tends to be lower than on the test dataset. Excluding LLEMMA and the appropriateness classifier’s precision on MAmmoTH outputs, each classifier has fairly high precision, recall and F1, suggesting they are effectively able to label solvable, accurate, and appropriate questions. The classifiers tend to have low ROC AUC across sources, suggesting they do not perform well at different prediction cutoffs. As a whole, Table 6 provides evidence that text classifiers can learn to label MATHWELL outputs, but future research should use more balanced datasets to improve their ability to label unsolvable, inaccurate, and inappropriate questions.
5 Conclusions
We explore context-free word problem generation and create MATHWELL, which generates a question/answer pair based only on an optional topic. To train our model, we generate synthetic data and use expert annotators to identify a high-quality training subset. We release SGSM, a synthetic dataset of 20,490 question/answer pairs for use in future research. Our evaluations show that MATHWELL outperforms other open-source LLMs at context-free word problem generation and that SGSM is of comparable quality to existing math QA data. These findings suggest that context-free word problem generation is a feasible and practical alternative to traditional context-dependent generators. Future research should train context-free word problem generators that can create questions aligned with specific math topics and grade levels.
Limitations
One important limitation of MATHWELL is that it is not designed to generate questions aligned with pre-specified grade levels and mathematical operations/topics, which we chose to leave to future research due to the high cost of annotating questions for these characteristics but would improve the model’s utility in classroom settings. Additionally, MATHWELL is trained and evaluated for generating word problems/solutions for K-8 students only; therefore, we do not recommend using it to generate question/answer pairs for other grade levels or for other tasks.
Another limitation of this work is the subjective nature of the appropriateness criteria. While it is critical model-generated questions are appropriate for students, it is hard to fully define all aspects of appropriateness and individuals may have differing opinions on the degree to which a question is appropriate or not. We chose to define several common reasons questions may not be appropriate for students (see Figure 9) and use annotators with K-12 teaching experience/training and who are familiar with what is appropriate in a school setting, but future research should continue to define this criteria and include multiple evaluators.
Ethics Statement
All data used to train MATHWELL come from open-access datasets and, therefore, should not contain any private sensitive information. MATHWELL may generate questions that are inappropriate for use in educational contexts and additional research should be conducted on the model before deploying it in classroom settings. Specifically, future research should continue to improve performance of text classifiers to filter out questions which are not appropriate for students.
Acknowledgements
The authors thank Zooniverse Zooniverse for providing a free and user-friendly platform for data annotation as well as our expert volunteer annotators for providing high-quality labels and feedback on our evaluation criteria/directions.
Appendix A Important Criteria for Context-free Word Problem Generator Training Data
The characteristics we find most important for training context-free word problem generators are PoT solutions written as Python functions and questions that are written at an appropriate mathematical difficulty for K-8 students. Regarding the former, when we modified our prompt to ask for a Python function solution instead of a Python code solution, the share of question/answer pairs with executable code from an early version of MATHWELL increased from 18.9% to 29.0%. For appropriateness, when we used the GSM-Hard Gao et al. (2023) dataset as part of MATHWELL’s training data, the model often generated questions with large numbers that are inappropriate for K-8 students. As shown in Table 1, the SGSM Train subset is the only math QA dataset that has these two characteristics.
Appendix B Annotated Data Characteristics
Our annotated dataset consists of 4,234 question/answer pairs with human annotations for solvability, accuracy, appropriateness, and MaC. The data are comprised of the 3,234 word problem dataset used to generate training data for MATHWELL in addition to the 250 evaluation set for each model described in Section 4. Based on our annotations, 81.1% of the question/answer pairs are solvable, 86.5% have accurate solutions, 67.3% are appropriate, and 57.4% meet all criteria.
Appendix C Additional Analyses
C.1 Logistic Regression for Predicting MaC
As shown in Table 7, the coefficients for all models for MaC remain negative and statistically significant relative to MATHWELL when controlling for the type of mathematical operation. This finding supports the assertion that MATHWELL is more capable of generating MaC questions regardless of the operation considered, even if it is less likely to generate questions from the more complex mathematical operations.
| Predictor | Coefficient | SE | Z | p |
| Constant | 1.648 | 0.182 | 9.053 | 0.000** |
| LLEMMA | -2.441 | 0.267 | -9.138 | 0.000** |
| MAmmoTH | -1.009 | 0.231 | -4.363 | 0.000** |
| Llama-2 | -0.587 | 0.241 | -2.435 | 0.015* |
| Constant | 1.237 | 0.286 | 4.327 | 0.000** |
| LLEMMA | -2.215 | 0.278 | -7.954 | 0.000** |
| MAmmoTH | -0.855 | 0.241 | -3.548 | 0.000** |
| Llama-2 | -0.492 | 0.245 | -2.006 | 0.045* |
| Addition | 0.187 | 0.195 | 0.956 | 0.339 |
| Subtraction | 0.439 | 0.207 | 2.124 | 0.034* |
| Multiplication | 0.215 | 0.212 | 1.014 | 0.311 |
| Division | -0.155 | 0.260 | -0.594 | 0.552 |
| Fractions | -0.347 | 0.361 | -0.960 | 0.337 |
| Decimals | -0.211 | 0.281 | -0.752 | 0.452 |
C.2 Accuracy by Question Type
As shown in Table 8, MATHWELL’s accuracy does not differ significantly or substantively by operation, as its accuracy remains above 90% for all operations, while the other models have a significant and/or substantive gap in their accuracy for the operation they perform best on relative to the operation they perform worst on. While MAmmoTH outperforms MATHWELL for addition, multiplication, and division, MATHWELL performs better in the other three operations and in overall accuracy.
| Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. |
| LLEMMA | 76.2 | 72.3 | 63.4 | 56.0 | 50.0 | 63.2 |
| MAmmoTH | 96.5 | 96.3 | 96.8 | 97.6 | 87.5 | 91.3 |
| Llama-2 | 89.3 | 91.1 | 87.5 | 80.0 | 82.4 | 75.0 |
| MATHWELL | 96.1 | 97.4 | 94.5 | 91.3 | 100.0 | 94.1 |
C.3 FKGL for All Versus MaC Questions
As shown in Figure 5, the FKGL distribution for all versus MaC questions does not significantly vary for all models except for LLEMMA, whose MaC questions tend to have lower FKGL than its average question.
C.4 Additional Automatic Evaluation Comparisons
As shown in Table 9, PPL does not significantly vary for all models except for LLEMMA when comparing all to MaC generations. In the table, we also compare the BERTScore between all and MaC questions from each model and show they are similar. While LLEMMA has a longer average token length for all its generations than MATHWELL, MATHWELL’s MaC average token length is the longest of the models considered.
| Model | PPL | M PPL | BF1 | M BF1 | A/M BF1 | G BF1 | AL | M AL | NDC |
| LLEMMA | 3.82 (0.10) | 3.14 (0.10) | 84.3 | 85.3 | 84.6 | 84.1 | 56.4 (1.44) | 50.9 (2.89) | 8.41 (0.09) |
| MAmmoTH | 2.76 (0.03) | 2.74 (0.04) | 86.0 | 86.4 | 86.1 | 84.7 | 45.9 (1.13) | 44.4 (1.15) | 8.25 (0.08) |
| Llama-2 | 2.52 (0.03) | 2.52 (0.04) | 85.5 | 85.8 | 85.6 | 84.3 | 51.6 (0.99) | 49.8 (1.19) | 8.20 (0.07) |
| MATHWELL | 2.44 (0.03) | 2.43 (0.03) | 85.5 | 85.7 | 85.6 | 84.2 | 54.7 (0.86) | 54.1 (0.97) | 8.23 (0.08) |
Appendix D Prompting Process
D.1 Standard Prompt
Our standard prompt for interacting with the MATHWELL model is, "Write a grade school math word problem about {topic} and Python function with a commented out step-by-step solution to solve the word problem." In this prompt, topic is an optional argument, which we randomly select from a list of topics relevant to K-8 students. We begin every prompt with a random selection of 8-shot examples from SGSM Train.
D.2 Suggested Prompting Strategies
We find MATHWELL is more likely to generate executable code when given a topic than when a topic is not specified. For example, when prompting our finetuned Llama-2 model before further training it on the SGSM Train data, we found the model generated executable code 63.1% of the time when given a topic, and only 32.3% of the time when a topic was not specified. As a result, for evaluating models in this paper, we provide them with a randomly selected topic, which also gives us the ability to assess their ability to effectively generate topic-specific word problems. Additionally, this evaluation strategy is aligned with how a teacher or student would use the model in practice, as they would want the generated questions to align with a particular topic. Qualitative evaluations of model generations also revealed that MATHWELL is more likely to generate executable code when the topic is more specific. For example, if their desired topic is superheroes, a user would have a higher likelihood of receiving a generation with executable code by prompting with a specific superhero (e.g., Superman) than leaving the topic general (e.g., superheroes).
Appendix E Example Generations
The sections below display randomly sampled generations from each model, including examples of both questions that do not meet evaluation criteria and those that do.
E.1 LLEMMA
E.1.1 Solvability
Positive Example
Steph Curry makes 2133 free throws for the year. He hits 342 more free throws than he misses. How many free throws does Steph Curry miss?
Negative Example
There are 5175 Pokémon available to battle in Pokémon Sword and Shield. There are 310 ground-type Pokémon and 182 water-type Pokémon. There are 610 Pokémon that are not ground nor water type. How many Pokémon are ground-type or water-type?
Negative Example Reasoning
The question contains conflicting information in that it defines the number of ground and water type Pokémon twice.
E.1.2 Accuracy
Positive Example
Naruto eats 4 bananas a day. How many bananas does he eat in a month if there are 30 days in a month?
Solution:
def solution():
#Naruto eats 4 bananas a day
bananas_per_day = 4
#There are 30 days in a month
days_in_month = 30
#The answer is
result = bananas_per_day * days_in_month
return result
Negative Example
12 cats eat 400 pounds of cat food every Saturday morning. 15 cats eat 500 pounds of cat food on Sunday morning. How many pounds of cat food are eaten in total?
Solution:
def solution():
#12 cats eat 400 pounds of cat food every
#Saturday morning
cat_food_satuday_morning = 12*400
#15 cats eat 500 pounds of cat food on Sunday
#morning
cat_food_sunday_morning = 15*500
#In total,
result = cat_food_satuday_morning +
cat_food_sunday_morning
return result
Negative Example Reasoning
The model multiplies the number of cats by the amount of cat food eaten, rather than adding the two sums of cat food together.
E.1.3 Appropriateness
Positive Example
Hello Kitty makes 18 bracelets in 4 hours. How many bracelets per hour does she make?
Negative Example
Batman caught a baddie with his trademark punch. Each punch knocks out 7 baddies. If Batman has thrown 60 punches, how many baddies has he knocked out?
Negative Example Reasoning
While this question is comical, it is not appropriate for a K-8 student because it involves physically harming another person.
E.2 MAmmoTH
E.2.1 Solvability
Positive Example
The 5 Power Rangers, Red, Blue, Black, Yellow, and Pink, are eating pizza. Red ate 31 slices, Blue ate 28 slices, Black ate 23 slices, Yellow ate 21 slices, and Pink ate 15 slices. How many slices of pizza did they eat in total?
Negative Example
In Fortnite, the player has 5 health points. When the player is hit by an enemy, they lose 2 health points. How many health points does the player have left?
Negative Example Reasoning
This question defines how many health points a player loses when they get hit by an enemy, but it does not define how many times the player got hit, so it is impossible to determine how many health points they have left.
E.2.2 Accuracy
Positive Example
Harry Potter and his friends have just finished their exams and are looking forward to a well-deserved break. They decide to go on a camping trip together. They have 120 Galleons between them. They spend 30 Galleons on food in the morning and 20 Galleons on food in the afternoon. They have 20 Galleons left. How many Galleons did they spend in the evening?
Solution:
def solution():
# total amount of Galleons they had
total_galleons = 120
# amount spent on food in the morning
morning_food = 30
# amount spent on food in the afternoon
afternoon_food = 20
# amount left
left_galleons = 20
# calculate the amount spent in the evening
evening_food = total_galleons - morning_food -
afternoon_food - left_galleons
return evening_food
print(solution())
222MAmmoTH is prone to ending its generations with print(solution()) because its PoT training data always ends with a print statement Yue et al. (2023).
Negative Example
Captain Marvel has 100 friends on Facebook. She has 40 more friends than the average number of friends her friends have. How many friends does the average friend of Captain Marvel have?
Solution:
def solution():
#Captain Marvel has 100 friends on Facebook
captain_marvel_friends = 100
#She has 40 more friends than the average
#number of friends her friends have
captain_marvel_friends_more = 40
#The average number of friends her friends have
average_friends = (captain_marvel_friends -
captain_marvel_friends_more) / 2
return average_friends
Negative Example Reasoning
The model calculates an average rather than subtract the average number of friends Captain Marvel’s friends have from her number of friends.
E.2.3 Appropriateness
Positive Example
Taylor Swift has 11 Grammys, 29 AMAs, 12 CMAs, 8 ACMs and 35 BMAs. How many awards has she won in total?
Negative Example
The soccer team has 32 players. Each player has 2 legs. How many legs does the team have?
Negative Example Reasoning
While this question is solvable, it is not appropriate because it is strange to ask how many legs a soccer team has.
E.3 Llama-2
E.3.1 Solvability
Positive Example
The Pittsburgh Steelers have 11 players on the field. 3 of the players are from the 10th grade, 6 are from the 11th grade, and the rest are from the 12th grade. How many players are from the 12th grade?
Negative Example
The cheerleaders at the school are planning a pep rally. They need to make 200 pom-poms. Each pom-pom requires 20 minutes to make. How many hours will it take to make 200 pom-poms?
Negative Example Reasoning
The question does not define how many cheerleaders there are, so it is impossible to determine how long it will take them to make the pom-poms.
E.3.2 Accuracy
Positive Example
Barbie has 100 pairs of shoes. She gave 20 pairs of shoes to Ken, 30 pairs of shoes to Skipper, and 40 pairs of shoes to Stacie. How many pairs of shoes does Barbie have left?
Solution:
def solution():
#Barbie started with 100 pairs of shoes
shoes_initial = 100
#She gave 20 pairs of shoes to Ken
shoes_ken = 20
#She gave 30 pairs of shoes to Skipper
shoes_skipper = 30
#She gave 40 pairs of shoes to Stacie
shoes_stacie = 40
#The answer is
result = shoes_initial - shoes_ken -
shoes_skipper - shoes_stacie
return result
Negative Example
A basketball team scored 120 points in a game. The team scored 30 points in the first quarter, 35 points in the second quarter, 20 points in the third quarter, and 35 points in the fourth quarter. How many points did the team score in the second half of the game?
Solution:
def solution():
#The team scored 120 points in a game
total_points = 120
#30 points in the first quarter
points_first_quarter = 30
#35 points in the second quarter
points_second_quarter = 35
#20 points in the third quarter
points_third_quarter = 20
#35 points in the fourth quarter
points_fourth_quarter = 35
#The answer is
result = points_second_quarter +
points_third_quarter + points_fourth_quarter
return result
Negative Example Reasoning
The model correctly defines the variables, but it incorrectly adds the points scored in the second quarter to the total for the points scored in the second half.
E.3.3 Appropriateness
Positive Example
Patrick Mahomes is a professional football player. He is 6 feet and 3 inches tall. 1 inch is 2.54 centimeters. How many centimeters tall is Patrick Mahomes?
Negative Example
A cat has 100 kittens. 20 of them are calico, 30 are tabby, and the rest are siamese. How many kittens are siamese?
Negative Example Reasoning
This question is not based in reality, as it is not possible for one cat to birth 100 kittens, nor is it possible for them to be different breeds.
E.4 MATHWELL
E.4.1 Solvability
Positive Example
A soccer team has 11 players on the field at a time. They have 22 more players on the bench. How many players are on the team in total?
Negative Example
Steph Curry, the NBA star, is 6.2 feet tall. He wears shoes that are 1.5 inches thick. How tall is he in inches? (1 foot is 12 inches)
Negative Example Reasoning
This question is not solvable because it is vague: we do not know if we are calculating his height in inches with or without shoes on.
E.4.2 Accuracy
Positive Example
In the game, Super Mario can stomp on 30 Goombas, 20 Koopas, and 10 Piranha Plants. How many enemies does he stomp on total?
Solution:
def solution():
#Super Mario can stomp on 30 Goombas
goombas = 30
#20 Koopas
koopas = 20
#and 10 Piranha Plants
piranha_plants = 10
#The answer is
result = goombas + koopas + piranha_plants
return result
Negative Example
A Minecraft player has 100000 blocks. 20000 of the blocks are dirt, 30000 of the blocks are stone, 20000 of the blocks are wood, and 30000 of the blocks are diamond. How many of the blocks are not dirt, stone, wood, or diamond?
Solution:
def solution():
#100000 blocks
blocks_total = 100000
#20000 of the blocks are dirt
blocks_dirt = 20000
#30000 of the blocks are stone
blocks_stone = 30000
#20000 of the blocks are wood
blocks_wood = 20000
#30000 of the blocks are diamond
blocks_diamond = 30000
#The answer is
result = blocks_total - blocks_dirt - blocks_stone
- blocks_wood
return result
Negative Example Reasoning
The model correctly defines the necessary variables, but does not subtract the number of diamond blocks from the total number of blocks.
E.4.3 Appropriateness
Positive Example
LeBron James has 12000 points. He is 4000 points away from the all-time scoring record. How many more points does he need to average per game for the next 20 games to break the record?
Negative Example
A field hockey team has 11 players. 3 of them are forwards, 3 of them are midfielders, 3 of them are defenders, and 2 of them are goalies. How many forwards are there?
Negative Example Reasoning
This question is inappropriate to give to a student because it does not require any mathematical operations to solve. It directly defines the number of forwards on the team.
Appendix F Annotation Process
F.1 Annotators
All annotators had previous K-12 teaching experience or training, including a research team member who annotated every question. We had three primary annotators who reviewed at least 200 questions each in addition to our research team member.
F.2 Inter-Annotator Agreement
For annotating synthetic data to train MATHWELL, 998 questions were annotated by two people and 232 were annotated by three people. Annotators agreed on solvability of the time, accuracy of the time, appropriateness of the time, all three labels of the time, and MaC of the time. The agreement rates for accuracy and solvability are higher than reported in recent human evaluation studies that analyze human preferences in LLM outputs, and the agreement rates for appropriateness and MaC are on par with these studies Bai et al. (2022); Ouyang et al. (2022); Stiennon et al. (2022); Ziegler et al. (2020). As a result, we feel confident in the quality of our labels.
F.3 Handling Annotator Disagreement
For annotating synthetic data to train MATHWELL, if the question was reviewed by two annotators and they disagreed on one of the criteria, we labeled the example as not having the desired criteria. If the question was reviewed by three annotators and there was a disagreement on one of the criteria, we assigned the label with the majority vote.
F.4 Validating Final Evaluation Labels
For annotating the 250 samples from each model for our experiments reported in Section 4, we randomized questions from each model and had them blindly reviewed by one of our highly trained annotators with K-12 teaching experience. To evaluate the quality of these labels, we had 285 randomly reviewed by one additional annotator and 60 random reviewed by two additional annotators. The annotators agreed on solvability of the time, accuracy of the time, appropriateness of the time, all three labels of the time, and MaC of the time. Similar to above, the agreement rates for accuracy and solvability are higher than reported in recent studies that explore human alignment of LLM outputs, and the agreement rates for appropriateness and MaC are on par with these studies Bai et al. (2022); Ouyang et al. (2022); Stiennon et al. (2022); Ziegler et al. (2020).
Additional analysis reveals that most annotator disagreement (81.6%) was due to the primary annotator being more conservative than the additional annotators by labeling questions as not having the desired criteria when the additional annotators rated them as having the desired criteria. As a result, we chose to use the labels from the primary annotator when reporting final results to be conservative, though we also found the results do not vary when switching labels based on annotator disagreement. Annotators were least likely to disagree on labels for MATHWELL outputs and our primary annotator was not more likely to rate MATHWELL outputs as having the desired criteria than the additional annotators. Taken together, this evidence suggests our final labels are highly accurate.
F.5 Annotation Interface
We used Zooniverse Zooniverse to collect our human annotation data. Figures 6, 7, 8 and 9 show the instructions each annotator was given for each of our evaluation criteria.
Appendix G Early MATHWELL Experimentation
In addition to training context-free question/answer pair generators, we also experiment with training context-free question generation models. Our theory is that if we could train a model to generate questions effectively, we could pass those questions to a math QA model to retrieve answers automatically. To test this theory, we finetune both Llama-2 and MAmmoTH as question generators using the same training data discussed in Section 3.1, except for excluding the solution for each question and modifying the standard prompt to ask the model to generate a question only. We then sample and evaluate 100 generations from each model. We find that MAmmoTH performs better than Llama-2 at this task, but neither model performs optimally. For example, only 19% of the MAmmoTH generations include the requested topic and 52.6% are solvable. Therefore, based on the results we report in Table 2, we conclude that it is more efficient to train a context-free question/answer pair generator than question generator.